# ReynoldsFlow: Exquisite Flow Estimation via Reynolds Transport Theorem

### v1

@@ Abstract
Optical flow is a fundamental technique for motion estimation, widely applied in video stabilization, interpolation, and object tracking. Recent advancements in artificial intelligence (AI) have enabled deep learning models to leverage optical flow as an important feature for motion analysis. However, traditional optical flow methods rely on restrictive assumptions, such as brightness constancy and slow motion constraints, limiting their effectiveness in complex scenes. Deep learning-based approaches require extensive training on large domain-specific datasets, making them computationally demanding. Furthermore, optical flow is typically visualized in the HSV color space, which introduces nonlinear distortions when converted to RGB and is highly sensitive to noise, degrading motion representation accuracy. These limitations inherently constrain the performance of downstream models, potentially hindering object tracking and motion analysis tasks. To address these challenges, we propose Reynolds flow, a novel training-free flow estimation inspired by the Reynolds transport theorem, offering a principled approach to modeling complex motion dynamics. Beyond the conventional HSV-based visualization, denoted ReynoldsFlow, we introduce an alternative representation, ReynoldsFlow+, designed to improve flow visualization. We evaluate ReynoldsFlow and ReynoldsFlow+ across three video-based benchmarks: tiny object detection on UAVDB, infrared object detection on Anti-UAV, and pose estimation on GolfDB. Experimental results demonstrate that networks trained with ReynoldsFlow+ achieve state-of-the-art (SOTA) performance, exhibiting improved robustness and efficiency across all tasks. The code is available at https://github.com/wish44165/ReynoldsFlow.
@@ 1 Introduction
Optical flow estimates pixel velocities by analyzing intensity changes between consecutive frames, playing a fundamental role in motion estimation [6, 44, 20, 47], object detection [4, 1, 16, 47, 3, 38], scene reconstruction [47, 29], and video stabilization [41, 45, 22]. Historically, these applications have relied on traditional methods such as Horn-Schunck [8] and Lucas-Kanade [25], which are based on assumptions of local brightness constancy and spatial smoothness. However, these assumptions often break down in the presence of occlusions, fine-scale motion, or complex backgrounds. More recently, deep learning-based models, such as FlowNet [10] and FlowNet2 [17], have been proposed to learn optical flow. Additionally, some of these models directly integrate the learned optical flow into their architecture to warp images as an intermediate step for tasks like video frame interpolation. Although these models achieve impressive results, they typically require extensive training on synthetic datasets such as FlyingChairs and FlyingThings3D [10] and often struggle to generalize to real-world scenarios without further fine-tuning. In addition, the general approach to visualizing optical flow often uses HSV-based visualization, which aligns well with human perception by offering better discrimination power through Hue and Saturation. This makes it ideal for tasks such as color tuning, scene segmentation, and computer graphics, as highlighted in [24]. However, while HSV-based optical flow visualization enhances color saliency, it provides limited value for tasks where motion dynamics and structural variations are more critical than chromatic differences.
To preserve the training-free nature, we propose Reynolds flow, a novel optical flow framework grounded in the Reynolds transport theorem. By reinterpreting optical flow as a transport phenomenon of the light field associated with rigid motion, Reynolds flow models a broader range of transport phenomena, including complex motion dynamics and varying lighting conditions. In addition to the conventional HSV-based visualization, ReynoldsFlow, we introduce an enhanced representation, ReynoldsFlow+, which integrates optical flow magnitude, Reynolds flow magnitude, and frame intensity. This approach effectively distinguishes objects in videos with complex backgrounds and varying ambient lighting conditions, addressing the perceptual ambiguities caused by non-uniform color interpretation in HSV space. As demonstrated in Fig. 1, traditional HSV-based optical flow fails to convey accurate motion features in challenging datasets [9, 18, 27], whereas ReynoldsFlow+ provides a more explicit perceptual representation of motion characteristics. Our contributions are as follows:
@@
We introduce an enhanced visualization, ReynoldsFlow+, improving motion clarity for downstream tasks.
@@ 5 Conclusion
This work proposes Reynolds flow, a training-free optical flow estimation framework grounded in the Reynolds transport theorem that generalizes traditional methods. Building on this framework, we introduce ReynoldsFlow and ReynoldsFlow+. We train YOLOv11n using these flow videos. Our experiments demonstrate that ReynoldsFlow+ provides more informative motion features even without the direction information of the flows. We evaluated these methods on diverse tasks, including object detection on UAVDB and Anti-UAV and pose estimation on GolfDB. ReynoldsFlow+ consistently improves accuracy on varying scales and complexities, serving as a practical preprocessing module that mitigates downstream learning difficulty without additional computational overhead. In addition, its high computational efficiency makes ReynoldsFlow+ ideal for real-time deployment on resource-constrained edge devices. Future work will explore the integration of ReynoldsFlow+ with learning-based systems on FPGAs to enhance scalability and versatility further.

### v2

@@ Abstract
Optical flow is a fundamental technique for motion estimation, widely applied in video stabilization, interpolation, and object tracking. Traditional optical flow estimation methods rely on restrictive assumptions like brightness constancy and slow motion constraints. Recent deep learning-based flow estimations require extensive training on large domain-specific datasets, making them computationally demanding. Also, artificial intelligence (AI) advances have enabled deep learning models to take advantage of optical flow as an important feature for object tracking and motion analysis. Since optical flow is commonly encoded in HSV for visualization, its conversion to RGB for neural network processing is nonlinear and may introduce perceptual distortions. These transformations amplify the sensitivity to estimation errors, potentially affecting the predictive accuracy of the networks. To address these challenges that are influential to the performance of downstream network models, we propose Reynolds flow, a novel training-free flow estimation inspired by the Reynolds transport theorem, offering a principled approach to modeling complex motion dynamics. In addition to conventional HSV-based visualization of Reynolds flow, we also introduce an RGB-encoded representation of Reynolds flow designed to improve flow visualization and feature enhancement for neural networks. We evaluated the effectiveness of Reynolds flow in video-based tasks. Experimental results on three benchmarks, tiny object detection on UAVDB, infrared object detection on Anti-UAV, and pose estimation on GolfDB, demonstrate that networks trained with RGB-encoded Reynolds flow achieve SOTA performance, exhibiting improved robustness and efficiency across all tasks. The code is available at https://github.com/wish44165/ReynoldsFlow.
@@ 1 Introduction
Optical flow estimates pixel velocities by analyzing intensity changes between consecutive frames, playing a fundamental role in motion estimation [6, 44, 20, 47], object detection [4, 1, 16, 47, 3, 38], scene reconstruction [47, 29], and video stabilization [41, 45, 22]. Historically, these applications have relied on traditional methods such as Horn-Schunck [8] and Lucas-Kanade [25], which are based on assumptions of local brightness constancy and spatial smoothness. However, these assumptions often break down in the presence of occlusions, fine-scale motion, or complex backgrounds. More recently, deep learning-based models, such as FlowNet [10] and FlowNet2 [17], have been proposed to learn optical flow. Additionally, some of these models directly integrate the learned optical flow into their architecture to warp images as an intermediate step for tasks like video frame interpolation. Although these models achieve impressive results, they typically require extensive training on synthetic datasets such as FlyingChairs and FlyingThings3D [10] and often struggle to generalize to real-world scenarios without further fine-tuning. In addition, the general approach to visualizing optical flow often uses HSV-based visualization, which has better discrimination power through Hue and Saturation. This makes it ideal for tasks such as color tuning, scene segmentation, and computer graphics, as highlighted in [24]. However, while HSV-based optical flow visualization enhances color saliency, nonlinear perceptional sensitivity of the transformation between HSV and RGB limited its value for tasks such as moving object tracking, where structural variations and varying environmental lighting are more critical than chromatic differences.
To preserve the training-free nature, we propose Reynolds flow, a novel optical flow framework grounded in the Reynolds transport theorem. By reinterpreting optical flow as a transport phenomenon of the light field associated with rigid motion, Reynolds flow models a broader range of transport phenomena under complex motion dynamics and varying lighting conditions. In addition to the conventional HSV-based visualization, denoted as ReynoldsFlow, we introduce an enhanced representation in RGB space, denoted as ReynoldsFlow+, which integrates magnitudes of image intensity, optical flow, and Reynolds flow. This approach effectively distinguishes moving objects in videos with complex backgrounds and varying ambient light, addressing the perceptual ambiguities caused by non-uniform color interpretation in HSV space. As demonstrated in Fig. 1, traditional HSV-based optical flow does not convey accurate motion characteristics in challenging datasets [9, 18, 27], while ReynoldsFlow+ provides a more explicit perceptual representation of motion characteristics. Our contributions are as follows:
@@
We introduce an enhanced visualization, ReynoldsFlow+, which improves motion clarity for downstream tasks.
@@ 5 Conclusion
This work proposes Reynolds flow, a training-free optical flow estimation framework grounded in the Reynolds transport theorem that generalizes traditional methods. Building on this framework, we introduce ReynoldsFlow and ReynoldsFlow+. Our experiments demonstrate that ReynoldsFlow+ provides more informative motion features even without the direction information of the flows. We evaluated these methods on diverse tasks, including object detection on UAVDB and Anti-UAV and pose estimation on GolfDB. ReynoldsFlow+ consistently improves accuracy on varying scales and complexities, serving as a practical preprocessing module that mitigates downstream learning difficulty without additional computational overhead. In addition, its high computational efficiency makes ReynoldsFlow+ ideal for real-time deployment on resource-constrained edge devices. Future work will explore the integration of ReynoldsFlow+ with learning-based systems on FPGAs to enhance scalability and versatility further.
6 Acknowledgments
This research is supported by the grants 112-2115-M-A49-011 and 113-2115-M-A49-002 from NSTC, Taiwan.